$SPAD/src/input f07fef.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
uplo:="L"
n:=4
nrhs:=2
a:Matrix SF:=
 [[2.04 ,0.0 ,0.0 ,0.0 ],_
  [-1.53 ,1.64 ,0.0 ,0.0 ],_
  [0.28 ,-0.25 ,0.79 ,0.0 ],_
  [-0.05 ,0.67 ,0.66 ,0.54 ]]
lda:=4
ldb:=4
b:Matrix SF:=
 [[8.7 ,8.3 ],_
  [-13.35 ,2.13 ],_
  [1.89 ,1.61 ],_
  [-4.14 ,5 ]]
result:=f07fef(uplo,n,nrhs,a,lda,ldb,b)

———-

References^{[1] nothing}

4